# FPT approximations for Capacitated Sum of Radii and Diameters

### v1

@@ 1 Introduction
We are interested in a much more challenging generalization where each point $p$ has an inherent capacity $U_{p}\geq 0$, indicating an upper bound on the number of points it can serve as a center of a cluster in the solution. [^4] Formally, following footnote 1, a valid solution to the capacitated version also contains an assignment $\sigma:P\rightarrow\{x_{1},\dots,x_{k}\}$ such that if $\sigma(p)=x_{i}$ then $\delta(p,x_{i})\leq r_{i}$, and for every $i$, $|\sigma^{-1}(x_{i})|\leq U_{x_{i}}$. This is known as the Non-Uniform Capacitated clustering problem. If all the points have the same capacity $U$, the problem is referred to as the Uniform Capacitated clustering problem. Capacitated clustering models various applications, e.g. load balancing (where each cluster representative can handle only bounded number of objects). Capacitated clustering were thoroughly studied: Capacitated $k$-center admits a constant factor approximation in polynomial time for both uniform [KS00] and non-uniform capacities [CHK12, ABC+15]. There are several bi-criteria polytime approximations for $k$-median [CR05, BFRS15, BRU16, DL16, Li16, Li17].
There are FPT approximation algorithms for capacitated $k$-median and capacitated $k$-means with approximation factors $(3+\varepsilon)$ and $(9+\varepsilon)$ respectively [CL19] (improving over [ABM+19, XZZ19]). From the other hand, assuming gap-ETH [^5] Informally, gap-ETH says that there exists an $\varepsilon>0$ such that no sub-exponential time algorithm for $3$-SAT can distinguish whether a given $3$-SAT formula has a satisfying assignment or every assignment satisfies at most $(1-\varepsilon)$ fraction of the clauses., in FPT  time it is impossible to approximate $k$-median and $k$-means with factors better than $(1+\frac{2}{e})$ and $(1+\frac{8}{e})$, respectively (even without capacities) [CGK+19]. There is a $4+\varepsilon$ FPT-approximation for uniform capacitated $k$-center [JKY24, GJ23] [^6] In fact, [GJ23] obtained a $2$-approximation for capacitated $k$-center under soft assignments (where different centers can be co-located). In contrast, in this paper we consider only hard assignments (where we can open only a single center at each point). One can use [GJ23] to obtain a $4$-approximation for uniform capacitated $k$-center w.r.t. hard assignments. See Section 1.3 for a further discussion..
@@
From the lower bound side, Jaiswal et. al. [JKY24] showed that assuming ETH [^7] Informally, ETH (Exponential Time Hypothesis), says that there is no subexponential time algorithm for $3$-SAT, there is some constant $\beta>0$, such that any $\beta$-approximation for CapSoR (with non-uniform capacities) requires $2^{\Omega(k/{\rm poly}\log k)}n^{O(1)}$ time. The first contribution of this paper is to show that CapSoR is APX hard even with uniform capacities. That is compared to [JKY24] we removed the ETH assumption, and used only uniform capacities. This is in contrast to the uncapacitated version which is belived to admit a PTAS (assuming $NP\not\subseteq QP$).
@@
There exists $\alpha>1$ such that it is NP-hard to approximate Uniform CapSoR better than factor $\alpha$.
@@
Assuming gap-ETH, there is a fixed $\alpha>1$ such that there is no $f(k)n^{o(k)}$ time $\alpha$-approximation algorithm for Uniform CapSoR.
Note that if we allow $n^{O(k)}$ time, then we can exactly solve Sum of Radii  using brute-force, even with non-uniform capacities (see footnote 13). Therefore, Theorem 2 not only rules out FPT-approximation algorithms for achieving an $\alpha$-approximation for this problem but also implies that in this context, the best algorithm for obtaining $\alpha$-approximation is the naive brute-force algorithm, which runs in $n^{O(k)}$ time.
@@
There is a (deterministic) FPT-approximation algorithm that finds a $(3+2\sqrt{2}+\varepsilon)$-approximation for Non-uniform CapSoR for any $\varepsilon>0$, and runs in time $2^{O(k^{2}+k\log(k/\varepsilon))}n^{O(1)}$. Furthermore, the algorithm yields $(3+2\sqrt{2}+\varepsilon)$-approximation for non-uniform capacities even when the objective is a monotone symmetric norm of the radii.
@@
Lower Bound $\ell_{1}$ $\beta>1$ $2^{\Omega(\frac{k}{{\rm polylog}k})}\cdot{\rm poly}(n)$ (ETH) [JKY24]
Uniform Lower Bound $\ell_{1}$ exact $n^{\Omega(k)}$ (ETH) [BLS23]
                              APX-hard Theorem 1
$\alpha>1$ $n^{\Omega(k)}$ (gap-ETH) Theorem 2
Non-Uniform Capacities $\ell_{1}$ $15+\varepsilon$ $2^{O(k^{2}\log(k/\varepsilon))}n^{O(1)}$  [BLS23]
@@
general $3+\sqrt{8}+\varepsilon\approx 5.83$ $2^{O(k^{2}+k\log(k/\varepsilon))}n^{O(1)}$ Theorem 3
@@
Non-Uniform general $7+\varepsilon$ $2^{O(k^{2}+k\log(k/\varepsilon))}n^{O(1)}$ Theorem 6
Table 1: Summary of our results and previous results for CapSoR and CapSoD. The results for CapSoD are implicit in the corresponding references. Furthermore, CapSoD makes sense only for the cluster capacities. While CapSoR with the $\ell_{\infty}$ objective corresponds to the Capacitated $k$-Center problem, CapSoD with the $\ell_{\infty}$ objective corresponds to the Capacitated Max-$k$-Diameter problem. The inapproximibility result, assuming gap-ETH, rules out any $f(k)n^{o(k)}$ time algorithm for uniform CapSoR ($\ell_{1}$ objective) that obtains an $\alpha$-approximation, for some fixed constant $\alpha$ bounded away from $1$.
@@ 1.1 Other Norm Objectives
In the Sum of Radii problem the goal is to choose $k$ balls covering all the metric points such that the sum of radii $\sum_{i}r_{i}$ is minimized. A natural generalization is to optimize some other objectives. Specifically, given a norm $\|\cdot\|:\mathbb{R}^{k}\rightarrow\mathbb{R}_{\geq 0}$, choose balls covering all metric points so that to minimize the norm of the radii vector $\|(r_{1},\dots,r_{k})\|$. One natural example is the $\ell_{p}$ norm where $\|(r_{1},\dots,r_{k})\|_{p}=\left(\sum_{i=1}^{k}r_{i}^{p}\right)^{1/p}$. Note that the $\ell_{1}$ norm objective is exactly the Sum of Radii problem, while the $\ell_{\infty}$ norm objective ( $\|(r_{1},\dots,r_{k})\|_{\infty}=\max_{i\in[1,k]}r_{i}$ ) is the $k$-center problem. The other $\ell_{p}$-norm objectives interpolate between these two fundamental problems in clustering. The power of such generalizations have been recently studied [CS19, ABB+23], which surprisingly not only unifies the existing FPT-AS for several different problems such as $k$-median, $k$-means, and $k$-center, but also yields FPT-AS for advanced problems such as priority $k$-center [GW06, BCCN21, Ple87], $\ell$-centurm [NP05, LNG15], ordered $k$-median[BSS18, BJKW19], and Socially Fair $k$-median [AGGN10, BCMN14, MV21, GSV21], which were previously unresolved.
@@
For non-uniform Capacitated Sum of Radii, Jaiswal et. al. [JKY24] obtained approximation factor of $(4+\sqrt{13}+\varepsilon)\approx 7.61$ w.r.t. any $\ell_{p}$-norm objective in $2^{O(k^{3}+k\log(k/\varepsilon))}n^{O(1)}$ time. As stated in Theorem 3 above, for non-uniform Capacitated Sum of Radii, we obtain approximation factor of $(3+\sqrt{8}+\varepsilon)\approx 5.83$ w.r.t. any monotone symmetric norm objective, in $2^{O(k^{2}+k\log(k/\varepsilon))}n^{O(1)}$ time. Thus we improve over [JKY24] on three fronts (see Table 1): (1) approximation factor, (2) generalizing to any monotone symmetric norm from $\ell_{p}$ norms, and (3) running time.
@@ 1.2 Sum of Diameters and cluster capacities
A closely related problem is the Sum of Diameters problem, which has been studied extensively and predates Sum of Radii [HJ87, HJ97, MS89, CRW91, DMR+00, CP04, BS15, DMR+00, FJ22]. Here, the goal is to partition the point set $P$ into $k$ clusters $C_{1},\dots,C_{k}$, and the objective is to minimize the sum of cluster diameters $\sum_{i=1}^{k}{\rm diam}(C_{i})$, where the diameter of a cluster is ${\rm diam}(C_{i})=\max_{x,y\in C_{i}}\delta(x,y)$ the maximum pairwise distance between two cluster points. Note that, unlike Sum of Radii, in this problem, there are no centers representing the clusters. Furthermore, this problem is NP-hard to approximate to a factor better than $2$ in polynomial time [DMR+00] (unlike Sum of Radii, which admits a QPTAS). A simple observation shows that any $\alpha$-factor approximation for Sum of Radii implies $2\alpha$-factor approximation for Sum of Diameters in a black-box way (and vice-versa). [^9] Given a set of points $P$, denote by $\mathcal{R}$ and $\mathcal{D}$ the value of the optimal solutions to Sum of Radii and Sum of Diameters respectively. It holds that $\mathcal{R}\leq\mathcal{D}\leq 2\cdot\mathcal{R}$. Indeed, consider an optimal solution to Sum of Diameters of cost $\mathcal{D}$, then by picking arbitrary center in each cluster we obtain a solution for Sum of Radii of cost at most $\mathcal{D}$ (thus $\mathcal{R}\leq\mathcal{D}$). From the other hand, given a solution to Sum of Radii of cost $\mathcal{R}$, the clusters induced by the balls constitute a solution to Sum of Diameters of cost at most $2\cdot\mathcal{R}$ (thus $\mathcal{D}\leq 2\cdot\mathcal{R}$). This trick has often been used to design approximation algorithms for Sum of Diameters. In fact, the current state-of-the-art algorithms for Sum of Diameters, including a polynomial-time $(6+\varepsilon)$-approximation [BERW24] and a quasi-polynomial-time $(2+\varepsilon)$-approximation [GKK+10], are based on this implicit trick by applying it to the polynomial-time $(3+\varepsilon)$-approximation and QPTAS for Sum of Radii, respectively.
@@
We introduce the problem of Capacitated Sum of Diameters (CapSoD). Here we are initially given $k$ capacities $U_{1},\dots,U_{k}$, the goal is to partition the point set $P$ into $k$ clusters $C_{1},\dots,C_{k}$, such that for every $i$, $|C_{i}|\leq U_{i}$, while the objective is to minimize the sum of cluster diameters. If all the capacities are equal $U_{1}=\dots=U_{k}$, we will call the capacities uniform (and otherwise as non-uniform). Under uniform capacities, the reduction mentioned in footnote 9 goes through. Thus an $\alpha$ approximation algorithm to uniform capacitated Sum of Radii transfers in a black-box manner into a $2\alpha$ approximation algorithm for the uniform capacitated Sum of Diameters, with the same running time. In particular, by using [JKY24], one can obtain a $6+\varepsilon$-approximation for uniform capacitated Sum of Diameters in FPT time. In fact, similarly to Sum of Radii, one can study Capacitated Sum of Diameters w.r.t. any norm objective, and the reduction will still go through. Thus it follows from [JKY24] that for any $\ell_{p}$ norm objective, uniform capacitated Sum of Diameters admits $2(1+\varepsilon)(4^{p-1}+1)^{1/p}$ FPTapproximation. Similarly, using our Theorem 4, we can obtain FPTtime $6+\varepsilon$ approximation for uniform capacitated Sum of Diameters w.r.t. to any monotone symmetric norm, or $2(1+\varepsilon)(3^{p-1}+1)^{1/p}$-approximation w.r.t. $\ell_{p}$ norm objective. Sum of Diameters is a fundamental and important problem. It’s capacitated version was not previously explicitly studied simply because there was nothing to say beyond this simple reduction.
@@
Bandyapadhyay et. al. [BLS23] introduced the problem of Capacitated Sum of Radii where each point $p\in P$ has a capacity $U_{p}$, and a cluster centered in $p$ can contain at most $U_{p}$ points. This corresponds for example to a scenario where we want to construct water wells, and a well constructed at point $p$ can serve up to $U_{p}$ clients. However, an equally natural problem is where one is given $k$ capacities $U_{1},\dots,U_{k}$, and the goal is to construct $k$ clusters with arbitrarily centers, such that the $i$’th cluster contains at most $U_{i}$ points. This problem is similar to our capacitated Sum of Diameters, and can correspond to a scenario where one want to distribute already existing $k$ water tanks (for example in a tent village during a festival). We refer to the two versions of the problem as node capacities, and cluster capacities, respectively. Note that for uniform capacities the two versions coincide. Further, note that the reduction from Sum of Radii to Sum of Diameters mentioned in footnote 9 holds in the capacitated version w.r.t. cluster capacities. Our results on node capacities in Theorem 3 hold for cluster capacities as well ($3+2\sqrt{2}+\varepsilon$ approximation for any monotone symmetric norm objective). No results on Sum of Radii with non-uniform cluster capacities were previously known.
1.3 Further Related Work
[…]

### v2

@@ 1 Introduction
We are interested in a much more challenging generalization where each point $p$ has an inherent capacity $U_{p}\geq 0$, indicating an upper bound on the number of points it can serve as a center of a cluster in the solution. [^4] Formally, following footnote 1, a valid solution to the capacitated version also contains an assignment $\sigma:P\rightarrow\{x_{1},\dots,x_{k}\}$ such that if $\sigma(p)=x_{i}$ then $\delta(p,x_{i})\leq r_{i}$, and for every $i$, $|\sigma^{-1}(x_{i})|\leq U_{x_{i}}$. This is known as the Non-Uniform Capacitated clustering problem. If all the points have the same capacity $U$, the problem is referred to as the Uniform Capacitated clustering problem. Capacitated clustering naturally models many applications: in load balancing, cluster centers correspond to servers with limited service capacity, the radius captures latency, and minimizing the sum of radii models reducing aggregate latency subject to serve load; in wireless sensors, each cluster head can accommodate only bounded number of devices due to bandwidth, memory, or battery limitations, the radius corresponds to communication range, and optimizing the sum of radii reflects minimizing total communication energy. Capacitated clustering is thoroughly studied: Capacitated $k$-center admits a constant factor approximation in polynomial time for both uniform [KS00] and non-uniform capacities [CHK12, ABC+15]. There are several bi-criteria polytime approximations for $k$-median [CR05, BFRS15, BRU16, DL16, Li16, Li17].
There are FPT approximation algorithms for capacitated $k$-median and capacitated $k$-means with approximation factors $(3+\varepsilon)$ and $(9+\varepsilon)$ respectively [CL19] (improving over [ABM+19, XZZ19]). From the other hand, assuming gap-ETH [^5] Informally, gap-ETH says that there exists an $\varepsilon>0$ such that no sub-exponential time algorithm for $3$-SAT can distinguish whether a given $3$-SAT formula has a satisfying assignment or every assignment satisfies at most $(1-\varepsilon)$ fraction of the clauses., in FPT  time it is impossible to approximate $k$-median and $k$-means with factors better than $(1+\frac{2}{e})$ and $(1+\frac{8}{e})$, respectively (even without capacities) [CGK+19]. There is a $4+\varepsilon$ FPT-approximation for uniform capacitated $k$-center [JKY24, GJ23] [^6] In fact, [GJ23] obtained a $2$-approximation for capacitated $k$-center under soft assignments (where different centers can be co-located). In contrast, in this paper we consider only hard assignments (where we can open only a single center at each point). One can use [GJ23] to obtain a $4$-approximation for uniform capacitated $k$-center w.r.t. hard assignments. See Section 3 for a further discussion..
@@
From the lower bound side, Jaiswal et. al. [JKY24] showed that assuming ETH [^7] Informally, ETH (Exponential Time Hypothesis), says that there is no subexponential time algorithm for $3$-SAT., there is some constant $\beta>0$, such that any $\beta$-approximation for CapSoR (with non-uniform capacities) requires $2^{\Omega(k/{\rm poly}\log k)}n^{O(1)}$ time. The first contribution of this paper is to show that CapSoR is APX hard for an explicit factor and even with uniform capacities. That is compared to [JKY24] we removed the ETH assumption, used only uniform capacities, and showed an explicit factor of inapproximability. This is in contrast to the uncapacitated version which is believed to admit a PTAS (assuming $NP\not\subseteq QP$).
@@
For every constant $\varepsilon>0$, it is NP-hard to approximate Uniform CapSoR to within a factor of $(1+1/e-\varepsilon)$.
@@
Assuming gap-ETH, for every $\varepsilon>0$ and any function $f$, there is no $f(k)n^{o(k)}$ time algorithm that approximates Uniform CapSoR to within a factor of $(1+1/e-\varepsilon)$.
Note that if we allow $n^{O(k)}$ time, then we can exactly solve Sum of Radii using brute-force, even with non-uniform capacities (see footnote 14). Therefore, Theorem 2 not only rules out FPT-approximation algorithms for achieving approximation better than factor of $(1+1/e-\varepsilon)$ for this problem but also implies that in this context, essentially the best algorithm is the naive brute-force algorithm, which runs in $n^{O(k)}$ time.
@@
There is a (deterministic) FPT-approximation algorithm that finds a $(3+2\sqrt{2}+\varepsilon)$-approximation for Non-uniform CapSoR for any $\varepsilon>0$, and runs in time $2^{O(k^{2}\log k+k\log(k/\varepsilon))}n^{O(1)}$. Furthermore, the algorithm yields $(3+2\sqrt{2}+\varepsilon)$-approximation for non-uniform capacities even when the objective is a monotone symmetric norm of the radii.
@@
(Non-Uni-)Hardness general $\beta>1$ $2^{\Omega(\frac{k}{{\rm polylog}k})}\cdot{\rm poly}(n)$ (ETH) [JKY24]
Hardness Uniform general exact $n^{\Omega(k)}$ (ETH) [BLS23]
NP-hard Theorem 1
$(1+1/e-\varepsilon)$ $n^{\Omega(k)}$ (gap-ETH) Theorem 2
Non-Uniform $\ell_{1}$ $15+\varepsilon$ $2^{O(k^{2}\log(k/\varepsilon))}n^{O(1)}$  [BLS23]
@@
general $3+\sqrt{8}+\varepsilon\approx 5.83$ $2^{O(k^{2}\log k+k\log(k/\varepsilon))}n^{O(1)}$ Theorem 3
@@
Non-Uniform general $7+\varepsilon$ $2^{O(k^{2}\log k+k\log(k/\varepsilon))}n^{O(1)}$ Theorem 6
Table 1: Summary of our results and previous results for CapSoR and CapSoD. The results for CapSoD are implicit in the corresponding references. Furthermore, CapSoD makes sense only for the cluster capacities. While CapSoR with the $\ell_{\infty}$ objective corresponds to the Capacitated $k$-Center problem, CapSoD with the $\ell_{\infty}$ objective corresponds to the Capacitated Max-$k$-Diameter problem. While our algorithms are deterministic for non-uniform capacities, the algorithms for the uniform capacity case are randomized. The inapproximability result, assuming gap-ETH, rules out any $f(k)n^{o(k)}$ time algorithm that can approximate uniform CapSoR (even with $\ell_{1}$ objective) to a factor better than $(1+1/e)$.
@@ 1.1 Other Norm Objectives
In the Sum of Radii problem the goal is to choose $k$ balls covering all the metric points such that the sum of radii $\sum_{i}r_{i}$ is minimized. A natural generalization is to optimize alternate objectives. Specifically, given a norm $\|\cdot\|:\mathbb{R}^{k}\rightarrow\mathbb{R}_{\geq 0}$, the task is to choose balls covering all metric points so as to minimize the norm of the radii vector $\|(r_{1},\dots,r_{k})\|$. A canonical example is the $\ell_{p}$ norm where $\|(r_{1},\dots,r_{k})\|_{p}=\left(\sum_{i=1}^{k}r_{i}^{p}\right)^{1/p}$. Note that the $\ell_{1}$ norm objective is exactly the Sum of Radii problem, while the $\ell_{\infty}$ norm objective ($\|(r_{1},\dots,r_{k})\|_{\infty}=\max_{i\in[1,k]}r_{i}$) is the $k$-center problem. The other $\ell_{p}$-norm objectives interpolate between these two fundamental problems in clustering. In particular, they capture objectives such as the sum of squared radii, [^9] Sum of squared radii is essentially equivalent to the $\ell_{2}$ objective: an $\alpha$-approximation algorithm for the $\ell_{2}$ objective yields an $\alpha^{2}$-approximation to the sum of squared radii. which has received attention in practice, especially in wireless network applications where it naturally models power consumption [AAB+06, LLX22]. The strength of such generalizations has been recently studied [CS19, ABB+23], which not only unified the existing FPT-AS for several different problems such as $k$-median, $k$-means, and $k$-center, but also lead to FPT-AS for advanced problems such as priority $k$-center [GW06, BCCN21, Ple87], $\ell$-centrum [NP05, LNG15], ordered $k$-median[BSS18, BJKW19], and Socially Fair $k$-median [AGGN10, BCMN14, MV21, GSV21], many of which were previously unresolved.
@@
For non-uniform Capacitated Sum of Radii, Jaiswal et. al. [JKY24] obtained approximation factor of $(4+\sqrt{13}+\varepsilon)\approx 7.61$ w.r.t. any $\ell_{p}$-norm objective in $2^{O(k^{3}+k\log(k/\varepsilon))}n^{O(1)}$ time. As stated in Theorem 3 above, for non-uniform Capacitated Sum of Radii, we obtain approximation factor of $(3+\sqrt{8}+\varepsilon)\approx 5.83$ w.r.t. any monotone symmetric norm objective, in $2^{O(k^{2}\log k+k\log(k/\varepsilon))}n^{O(1)}$ time. Thus we improve over [JKY24] on three fronts (see Table 1): (1) approximation factor, (2) generalizing to any monotone symmetric norm from $\ell_{p}$ norms, and (3) running time.
@@ 1.2 Sum of Diameters and Cluster Capacities
A closely related problem is the Sum of Diameters problem, which has been studied extensively and predates Sum of Radii [HJ87, HJ97, MS89, CRW91, DMR+00, CP04, BS15, DMR+00, FJ22]. Here, the goal is to partition the point set $P$ into $k$ clusters $C_{1},\dots,C_{k}$, and the objective is to minimize the sum of cluster diameters $\sum_{i=1}^{k}{\rm diam}(C_{i})$, where the diameter of a cluster is ${\rm diam}(C_{i})=\max_{x,y\in C_{i}}\delta(x,y)$ the maximum pairwise distance between two cluster points. Note that, unlike Sum of Radii, in this problem, there are no centers representing the clusters. Furthermore, this problem is NP-hard to approximate to a factor better than $2$ in polynomial time [DMR+00] (unlike Sum of Radii, which admits a QPTAS). A simple observation shows that any $\alpha$-factor approximation for Sum of Radii implies $2\alpha$-factor approximation for Sum of Diameters in a black-box way (and vice-versa). [^10] Given a set of points $P$, denote by $\mathcal{R}$ and $\mathcal{D}$ the value of the optimal solutions to Sum of Radii and Sum of Diameters respectively. It holds that $\mathcal{R}\leq\mathcal{D}\leq 2\cdot\mathcal{R}$. Indeed, consider an optimal solution to Sum of Diameters of cost $\mathcal{D}$, then by picking arbitrary center in each cluster we obtain a solution for Sum of Radii of cost at most $\mathcal{D}$ (thus $\mathcal{R}\leq\mathcal{D}$). On the other hand, given a solution to Sum of Radii of cost $\mathcal{R}$, the clusters induced by the balls constitute a solution to Sum of Diameters of cost at most $2\cdot\mathcal{R}$ (thus $\mathcal{D}\leq 2\cdot\mathcal{R}$). This trick has often been used to design approximation algorithms for Sum of Diameters. In fact, the current state-of-the-art algorithms for Sum of Diameters, including a polynomial-time $(6+\varepsilon)$-approximation [BERW24] and a quasi-polynomial-time $(2+\varepsilon)$-approximation [GKK+10], are based on this implicit trick by applying it to the polynomial-time $(3+\varepsilon)$-approximation and QPTAS for Sum of Radii, respectively.
@@
We introduce the problem of Capacitated Sum of Diameters (CapSoD). Here we are initially given $k$ capacities $U_{1},\dots,U_{k}$, the goal is to partition the point set $P$ into $k$ clusters $C_{1},\dots,C_{k}$, such that for every $i$, $|C_{i}|\leq U_{i}$, while the objective is to minimize the sum of cluster diameters. If all the capacities are equal $U_{1}=\dots=U_{k}$, we will call the capacities uniform (and otherwise as non-uniform). Under uniform capacities, the reduction mentioned in footnote 10 goes through. Thus an $\alpha$ approximation algorithm to uniform capacitated Sum of Radii transfers in a black-box manner into a $2\alpha$ approximation algorithm for the uniform capacitated Sum of Diameters, with the same running time. In particular, by using [JKY24], one can obtain a $6+\varepsilon$-approximation for uniform capacitated Sum of Diameters in FPT time. In fact, similarly to Sum of Radii, one can study Capacitated Sum of Diameters w.r.t. any norm objective, and the reduction will still go through. Thus it follows from [JKY24] that for any $\ell_{p}$ norm objective, uniform capacitated Sum of Diameters admits $2(1+\varepsilon)(4^{p-1}+1)^{1/p}$ FPT  approximation. Similarly, using our Theorem 4, we can obtain FPTtime $6+\varepsilon$ approximation for uniform capacitated Sum of Diameters w.r.t. to any monotone symmetric norm, or $2(1+\varepsilon)(3^{p-1}+1)^{1/p}$-approximation w.r.t. $\ell_{p}$ norm objective. Sum of Diameters is a fundamental and important problem. Its capacitated version was not previously explicitly studied simply because there was nothing to say beyond this simple reduction.
@@
Bandyapadhyay et. al. [BLS23] introduced the problem of Capacitated Sum of Radii where each point $p\in P$ has a capacity $U_{p}$, and a cluster centered in $p$ can contain at most $U_{p}$ points. This corresponds for example to a scenario where we want to construct water wells, and a well constructed at point $p$ can serve up to $U_{p}$ clients. However, an equally natural problem is where one is given $k$ capacities $U_{1},\dots,U_{k}$, and the goal is to construct $k$ clusters with arbitrary centers, such that the $i$’th cluster contains at most $U_{i}$ points. This problem is similar to our capacitated Sum of Diameters, and can correspond to a scenario where one wants to distribute already existing $k$ water tanks (for example in a tent village during a festival). We refer to the two versions of the problem as node capacities, and cluster capacities, respectively. Note that for uniform capacities the two versions coincide. Further, note that the reduction from Sum of Radii to Sum of Diameters mentioned in footnote 10 holds in the capacitated version w.r.t. cluster capacities. Our results on node capacities in Theorem 3 hold for cluster capacities as well ($3+2\sqrt{2}+\varepsilon$ approximation for any monotone symmetric norm objective). No results on Sum of Radii with non-uniform cluster capacities were previously known.